OpenStudy (anonymous):

did I do it right? [(1/2)-(2x^2)]^2 = (1/4)-(4x^4)-(2x^2)

OpenStudy (anonymous):

\[\left( a-b \right)^2=a^2+b^2-2ab\]

OpenStudy (anonymous):

You're very close, but one sign is wrong. Can you find it based on the formula just given?

OpenStudy (anonymous):

Yup, so it's actually, 1/4 + 4x^4 - 2x^2 ?

OpenStudy (anonymous):

Yes!

OpenStudy (anonymous):

Thanks!